when we subtract 10 from the distance between half a real number and -4 we get 4. find the number(s). show all work!
Answer by VFBundy(438)

(0.5n - (-4)) - 10 = 4

(0.5n + 4) - 10 = 4

0.5n - 6 = 4

0.5n = 10

n = 20

The number is 20.

===========================================

Check it:

Half of 20 is 10.

The distance between 10 and -4 is 14.

14 minus 10 is 4.

===========================================

You can also set it up this way:

(-4 - 0.5n) - 10 = 4

-4 - 0.5n - 10 = 4

-14 - 0.5n = 4

-0.5n = 18

n = -36

The number is -36.

============================================

Check it:

Half of -36 is -18.

The distance between -18 and -4 is 14.

14 minus 10 is 4.

=============================================

So...there are two answers...20 and -36.